15+35/5*8+12-7*4/2-6+
Answer by LinnW(1048)

15+35/5*8+12-7*4/2-6+
The first operations involve multiplication
and division.
for 35/5*8 , the division takes place first
since we want to go left to right with operations
at the same level.
This will produce 7*8
which = 56.
*****************
Now consider 7*4/2. Here multiplication takes
place first since we need to go left to right
7*4/2 = 28/2 = 14
*****************
Now back to the original expression,
15+35/5*8+12-7*4/2-6+
15+56+12-14-6
Now going left to right
71+12-14-6
83-14-6
69-6
63